OpenStudy (mellamokatie):

7. 16x^5 * 3x^7 A) 19x^12 B) 48x^35 C) 48x^12 D) 32x^12

OpenStudy (will.h):

Multiply the coefficients 16 * 3 = 48 Now multiply the variables with exponents when multiplying the variables with exponents we add the exponents So x^5 * x^7 = x^(5 + 7) So x^12 So the final answer is 48x^12